OpenStudy (anonymous):

5. Complete the following ordered pairs for the equation. 9x + -5y = 60 (-5, _) (_, -3) (10, _)

OpenStudy (ohmaigosshh):

First one... (-5, ) 9x + -5y = 60 9(-5) - 5y = 60 -45 - 5y = 60 Add 45 to both sides -5y = 105 Divide by -5 y = -21 .... so it will be (-5, -21)

OpenStudy (ohmaigosshh):

If you need help with the rest, just mention me :)